# Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery

### v1

@@ 6.1 Implications of Findings
  1. 1.
    Bridging Accuracy-Interpretability: RSAE shows a path to competitive predictive performance with transparency via sparse latent factors, vital in finance [83].
  2. 2.
    Enhanced Decision Support: Interpretable factors (e.g., ”Global Demand”) can aid hedging, inventory, and investment decisions.
  3. 3.
    Automated Factor Discovery: RSAE offers data-driven discovery of latent drivers and potential regime shift indicators [5].
  4. 4.
    Potentially Improved Risk Management: Validated sparse factors could enhance factor-based VaR or stress testing.
  5. 5.
    Broader Applicability: The architecture may apply to other financial/economic forecasting.
@@ 6.2 Limitations of the Study
    Interpretability Validation: Factor interpretation is qualitative and post-hoc; rigorous quantitative validation (event studies, causality tests) is needed [63].
@@
    Input Feature Dependence: Model learning is constrained by input feature quality and comprehensiveness.
@@
    Model Complexity/Cost: RSAE’s multi-component structure and inference loop increase training time and tuning effort.
@@
    Latent Variable Inference: Fixed $K$-step inference is an approximation; convergence and optimality are not guaranteed across all scenarios. Adaptive or convergence-based criteria for $\bm{z}^{*}$ inference are important future considerations.
@@
    Structural Breaks/Non-Stationarity: Adaptability to unforeseen structural breaks not in training data is a challenge [82].
@@
    Sparsity Level Selection: Choosing $\lambda$ involves a trade-off between interpretability and accuracy.
@@
    No Intrinsic Uncertainty Quantification: The model provides point forecasts, lacking probabilistic output [51].
@@
    Experimental Design: A single chronological train/validation/test split was used; more robust time-series cross-validation (e.g., rolling-origin or blocked CV) would provide more reliable out-of-sample performance estimates. Baselines received standard tuning (e.g., ‘$auto\_arima$‘, single CV for XGBoost), while RSAE had a specific grid search for $\lambda$; a fully balanced hyperparameter tuning budget across all models would strengthen comparisons. The directional accuracy metric employed is basic; more advanced metrics like the Pesaran-Timmermann test or Matthews Correlation Coefficient should be considered. The absence of statistical significance tests (e.g., Diebold-Mariano) for performance differences is a limitation.
@@ 6.3 Avenues for Future Research
    Advanced Interpretability/Validation: Integrate SHAP [66] or LIME [81]; use NLP on news to label/validate factors.
@@
    Improved Inference: Explore variational inference [55, 80] or adaptive optimization for $\bm{z}^{*}$.
@@
    Dynamic Sparsity Control: Allow $\lambda$ to be learned or vary with market context.
@@
    Regime-Switching Integration: Combine RSAE with Markov Switching Models [41].
@@
    Uncertainty Quantification: Extend to probabilistic forecasts (Bayesian NNs [71, 67], MC dropout [32]).
@@
    Hierarchical/Cross-Asset Modeling: Model related assets jointly to find shared drivers.
@@
    End-to-End Differentiable Optimization: Replace K-step inference with differentiable optimization layers [3, 39].
@@
    High-Frequency/Online Learning: Adapt for intraday data and online updates.
@@
    Reproducibility: Release code, preprocessed data, and random seeds to facilitate replication and extension.
@@
    Advanced Benchmarking: Compare against latest Transformer architectures (e.g., PatchTST) and apply rigorous statistical tests for performance differences (e.g., Diebold-Mariano tests, Model Confidence Sets). Evaluate with alternative directional accuracy metrics (e.g., Pesaran-Timmermann test).
@@ 6.4 Conclusion
This paper introduced the RSAE framework, a deep learning approach combining established techniques for multi-horizon commodity futures forecasting and interpretable latent factor discovery. By integrating time-lagged prediction with L1-regularized sparse inference, RSAE aims to balance predictive accuracy with model transparency. Empirical evaluation on Copper and Crude Oil futures data showed RSAE’s potential for competitive accuracy and for yielding sparse latent factors that correlate qualitatively with known market drivers. This offers an advantage over opaque black-box models. While acknowledging current limitations, including the qualitative nature of factor interpretation and the need for more extensive experimental validation (e.g., robust CV, statistical significance testing, broader hyperparameter tuning for baselines), the RSAE framework represents a promising direction for developing more insightful forecasting systems in finance. Future work focusing on advanced inference, uncertainty quantification, and rigorous validation is key to realizing its full potential.

### v2

@@ 6.1 Implications of Findings
RSAE offers a path to accuracy with transparency [86], enhanced decision support, automated (but to-be-validated) factor discovery [5], potential risk management tools, and broader applicability.
@@ 6.2 Limitations of the Study
    Interpretability validation is qualitative; quantitative methods are needed [65].
@@
    Input feature dependence constrains learning.
@@
    Model complexity implies higher computational cost/tuning effort.
@@
    The fixed $K=10$ inference steps were based on preliminary validation; convergence diagnostics (e.g., energy vs. iteration plots) and adaptive criteria are future work.
@@
    Adaptability to major structural breaks is a challenge [85].
@@
    Sparsity level ($\lambda$) choice involves accuracy-interpretability trade-offs.
@@
    Lacks intrinsic uncertainty quantification [53].
@@
    Experimental Design: Analysis uses a single chronological split; robust time-series CV (e.g., rolling-origin [10]) is needed. Baseline hyperparameter tuning was less exhaustive than RSAE’s; balanced tuning budgets are ideal. The DA metric is basic; alternatives (Pesaran-Timmermann test [79], MCC) are warranted. Formal statistical significance tests (Diebold-Mariano [26], MCS [44]) are crucial for validating performance differences.
@@ 6.3 Avenues for Future Research
    Advanced Interpretability: SHAP [68], LIME [84], NLP for factor labeling. SHAP on decoder for $\bm{z}^{*}\rightarrow\tilde{Y}$ contributions.
@@
    Improved Inference: Variational methods [57, 83], adaptive optimization.
@@
    Dynamic Sparsity Control: Learnable or context-varying $\lambda$.
@@
    Regime-Switching Integration [42].
@@
    Uncertainty Quantification: Bayesian NNs [73, 69], MC dropout [33].
@@
    Hierarchical/Cross-Asset Modeling.
@@
    End-to-End Differentiable Optimization [3, 40].
@@
    High-Frequency/Online Learning.
@@
    Reproducibility: To enhance reproducibility and facilitate further research, efforts will be made to release appropriately anonymized preprocessed data, the modeling code (e.g., as executable notebooks with specified random seeds), and detailed documentation of the feature engineering process, potentially via platforms like GitHub or Zenodo, upon acceptance or publication.
@@
    Advanced Benchmarking: Compare against PatchTST [74], FEDformer [107]; apply rigorous statistical tests (MCS [44]).
@@ 6.4 Conclusion
The RSAE framework, combining established deep learning techniques, shows potential for accurate, interpretable multi-horizon commodity futures forecasting. Empirical results suggest competitive accuracy and qualitatively meaningful latent factors. While acknowledging limitations (e.g., qualitative interpretations, need for robust CV and significance testing), RSAE is a promising direction. Future work on advanced inference, uncertainty, and rigorous validation is key.
Appendix A Latent Factor Correlation Analysis
This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in Section 5.3.2. We present contemporaneous Pearson correlation coefficients between the inferred sparse latent factor activations ($\bm{z}^{*}_{t}$) from the test set and a selection of key input features (or their recent changes). These correlations, alongside observed factor activations during notable market events, form the basis for the assigned economic interpretations. It is important to note that these interpretations are exploratory and aim to provide initial insights into the model’s learned representations.
A.1 Copper Latent Factor Correlations
For Copper, we highlight correlations for four prominent latent factors ($z^{*}_{Cu,1}$ to $z^{*}_{Cu,4}$) with selected macroeconomic and market-specific features. Table 5 shows the correlation coefficients, and Figure 3 visualizes these relationships.
Table 5: Pearson Correlation Coefficients: Copper Latent Factors vs. Selected Features (Test Set).
Feature $z^{*}_{Cu,1}$ (Demand/Inv.) $z^{*}_{Cu,2}$ (Macro/USD) $z^{*}_{Cu,3}$ (Supply Shock) $z^{*}_{Cu,4}$ (Sentiment)
Global Mfg. PMI (Change) $0.75$ $-0.10$ $0.05$ $-0.20$
LME Inventories (Change) $-0.65$ $0.05$ $0.15$ $0.10$
USD Index (DXY) (Change) $-0.25$ $0.70$ $-0.10$ $0.30$
Supply Disruption Index (Proxy) $0.10$ $0.00$ $-0.80$ $0.15$
VIX (Market Volatility) $-0.15$ $0.35$ $0.55$ $0.65$
Figure 3: Heatmap of Correlations: Copper Latent Factors vs. Selected Features (Test Set).
Interpretation of Copper Latent Factors:
  • Factor $z^{*}_{Cu,1}$ (Global Industrial Demand & Inventory Dynamics): This factor shows a strong positive correlation (+0.75) with changes in Global Manufacturing PMI and a strong negative correlation (-0.65) with changes in LME Copper Inventories. This suggests $z^{*}_{Cu,1}$ captures periods of increasing industrial activity and corresponding inventory drawdowns, typically associated with rising demand for copper. Its weaker negative correlation with the USD Index (-0.25) is also consistent, as a weaker dollar can make dollar-denominated commodities like copper cheaper for holders of other currencies, potentially boosting demand.
  • Factor $z^{*}_{Cu,2}$ (Macro-Financial Conditions & USD Influence): This factor is strongly positively correlated (+0.70) with changes in the USD Index (DXY). It also shows a moderate positive correlation (+0.35) with the VIX. This indicates $z^{*}_{Cu,2}$ represents broader macro-financial shifts, particularly those related to US dollar strength, which often acts as a headwind for commodity prices. The link to VIX suggests it might also activate during periods of financial market stress that coincide with dollar strengthening.
  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with a (hypothetical) Supply Disruption Index (where lower values of the index indicate more severe disruptions). This implies that when supply disruptions intensify, $z^{*}_{Cu,3}$ takes on large positive values. Its moderate positive correlation with VIX (+0.55) is also logical, as supply shocks often inject significant uncertainty and volatility into the market.
  • Factor $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): This factor exhibits a strong positive correlation (+0.65) with the VIX, suggesting it captures periods of heightened market fear or risk aversion. Its moderate positive correlation with the DXY change (+0.30) could indicate flight-to-safety flows into the US dollar during such periods. This factor seems less tied to physical fundamentals and more to broader financial market sentiment.
A.2 Crude Oil Latent Factor Correlations
Similarly, for Crude Oil, four key latent factors ($z^{*}_{Oil,1}$ to $z^{*}_{Oil,4}$) were analyzed against relevant features. Table 6 presents the correlation coefficients, and Figure 4 provides the heatmap.
Table 6: Pearson Correlation Coefficients: Crude Oil Latent Factors vs. Selected Features (Test Set).
Feature $z^{*}_{Oil,1}$ (US Demand/Inv.) $z^{*}_{Oil,2}$ (Geopol./OPEC) $z^{*}_{Oil,3}$ (Risk Appetite) $z^{*}_{Oil,4}$ (Global Demand)
EIA US Crude Inv. (Change) $-0.70$ $0.10$ $0.05$ $-0.15$
US Gasoline Demand (Proxy Change) $0.60$ $0.05$ $0.10$ $0.20$
Geopol./OPEC+ Event (Proxy) $0.15$ $0.80$ $0.30$ $0.00$
OVX (Oil Volatility) $-0.10$ $0.45$ $0.75$ $-0.25$
Global Growth Lead Ind. (Change) $0.20$ $0.10$ $-0.15$ $0.70$
Figure 4: Heatmap of Correlations: Crude Oil Latent Factors vs. Selected Features (Test Set).
Interpretation of Crude Oil Latent Factors:
  • Factor $z^{*}_{Oil,1}$ (US Demand, Inventory, and Refining Dynamics): This factor shows a strong negative correlation (-0.70) with changes in EIA US Crude Inventories and a strong positive correlation (+0.60) with a proxy for US Gasoline Demand. This clearly links $z^{*}_{Oil,1}$ to the state of the US oil market balance: higher values of this factor correspond to falling inventories and/or rising demand, typically bullish for WTI prices.
  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with a proxy for Geopolitical/OPEC+ supply-tightening events (where a positive value for the proxy indicates an event expected to reduce supply or increase risk premium). It also correlates moderately (+0.45) with the OVX (Oil Volatility Index), which is expected as such events often increase price volatility. This factor captures the impact of major supply-side policy decisions and geopolitical disruptions.
  • Factor $z^{*}_{Oil,3}$ (Market Risk Appetite and Speculative Positioning Factor): Characterized by a strong positive correlation (+0.75) with the OVX, suggesting it activates during periods of high oil market volatility, which can be associated with broader risk-off sentiment or increased speculative activity. Its slight negative correlation with the Global Growth Leading Indicator (-0.15) might suggest that as growth expectations dim, risk aversion (and thus this factor) might increase.
  • Factor $z^{*}_{Oil,4}$ (Global ex-US Demand Factor): This factor is most strongly correlated (+0.70) with changes in a Global Economic Growth Leading Indicator. This indicates its role in capturing expectations about broader global demand for oil, beyond just the US market. Its moderate positive correlation with the US Gasoline Demand proxy (+0.20) is also consistent with general demand trends.
These interpretations, based on correlations, provide a plausible understanding of how the RSAE model might be decomposing market information into sparse, meaningful drivers. Further validation using more formal attribution methods (as discussed in Section 6.3) would be required to confirm these relationships more rigorously.